# A Particle Algorithm for Mean-Field Variational Inference

### v2

@@ 7 Discussions
In this paper, we introduce PArticle VI (PAVI), a particle algorithm for mean-field variational inference (MFVI). The algorithm is based on formulating the relevant optimization problem in terms of Fokker-Planck equations and corresponding diffusion processes. We also provide a theoretical analysis of the particle algorithm, which provides a non-asymptotic upper bound on the $2$-Wasserstein distance between the empirical measure of the particles at the $n$th iteration and the solution of the MFVI optimization problem. Several future directions are worth pursuing. One may consider relaxing the regularity assumptions in Theorem 4.1 and adapting PAVI to the setting of parametric MFVI. Another interesting direction is to develop accelerated versions of PAVI using the recipe in the recent works Wang and Li (2022); Chen et al. (2023).

### v3

@@ 7 Discussions
We introduced PArticle VI (PAVI), a particle algorithm for mean-field variational inference (MFVI). The algorithm is based on formulating the relevant optimization problem in terms of Fokker-Planck equations and corresponding diffusion processes. Our theoretical analysis provides non-asymptotic $\mathcal{W}_{2}$ error bounds for the algorithm. Several future directions are worth pursuing. For instance, one may consider relaxing the regularity assumptions in Theorem˜4.1 and adapting PAVI to the setting of parametric MFVI. Another interesting direction is to develop accelerated versions of PAVI using the recipe in the recent works (wang2022accelerated; chen2023accelerating).